# Vision Transformers Don’t Need Trained Registers

### v4

@@ 5 Experiments
We evaluate how adding test-time registers affects the downstream performance of models trained without registers. We begin by detailing the evaluated models. Next, we compare performance across classification, dense prediction, unsupervised object discovery, and zero-shot segmentation tasks, finding that test-time registers perform comparably to their retrained counterparts. We then apply test-time registers to an off-the-shelf VLM to improve its interpretability while maintaining performance. Finally, we demonstrate register neurons’ ability to mask local information, moving their activations to adversarial image patches to prevent typographic attacks.
Models. We evaluate using DINOv2 (Oquab et al., 2024) and OpenCLIP (Cherti et al., 2023). For DINOv2, we use the publicly released ViT-L/14 checkpoints trained on LVD-142M, including both the standard model and a variant trained with four registers. These two models serve as our baselines, while our approach applies edits to the standard model. For OpenCLIP, we evaluate the ViT-L/14 and ViT-B/16 models trained on LAION-2B (Schuhmann et al., 2022). As no checkpoints with trained registers are available for OpenCLIP, we only compare our approach to the standard models.
@@
DINOv2 86.3 48.2 0.387
DINOv2 + trained registers 86.7 49.1 0.382
DINOv2 + test-time registers 86.3 49.1 0.378
OpenCLIP 77.4 40.2 0.603
OpenCLIP + test-time registers 77.5 40.3 0.595
@@
+ test-time registers 76.4
@@
+ test-time registers 71.3
@@
5.1 Classification and dense prediction
The results, presented in Table 3, show that models edited with test-time registers maintain their performance on ImageNet classification with slight performance gains over the base model on segmentation and depth estimation tasks. These improvements are consistent with the performance gains observed in DINOv2 explicitly trained with registers. Thus, we demonstrate that intervening on a model with test-time registers does not degrade the model’s representations, and in some cases, even enhances performance on prediction tasks. While we observe a small improvement in classification performance with the DINOv2 model trained with registers, we note that this was an independently trained model. Hence, the difference could be due to variations in initialization and training dynamics, rather than the effect of trained registers fixing artifacts. We present additional experiment details in Section A.8.
Zero-shot classification. We evaluate zero-shot ImageNet classification with OpenCLIP to assess whether test-time registers preserve the semantic structure of the original representation space. Unlike linear probing, where a trained linear head can compensate for small shifts in representation, zero-shot classification is more sensitive to such changes. We compare zero-shot performance before and after applying test-time registers across both ViT-L/14 and ViT-B/16 in Table 3, observing that the intervention does not sacrifice performance.
@@ 5.2 Zero-shot segmentation
To validate that test-time registers result in more interpretable attention maps, we follow a standard protocol for evaluating heatmap-based explainability methods (Hooker et al., 2019) – binarizing the heatmap into a foreground/background segmentation map, and evaluating its segmentation quality. We compute the mean attention map for the CLS token in the last layer (Chefer et al., 2021) for the original model without registers, the model with test-time registers, and a model trained with registers if available. We evaluate the zero-shot segmentation quality on ImageNet-segmentation (Guillaumin et al., 2014), which contains a subset of 4,276 images from the ImageNet validation set with annotated segmentations.
Results. We present our zero-shot segmentation scores in Table 4. We also qualitatively compare the attention maps for DINOv2 between test-time and trained registers in Figure 6, demonstrating similarly high-quality attention maps. Using test-time registers on both DINOv2 and OpenCLIP outperforms the original model on mean IOU and mAP with minimal drops in pixel accuracy. We suspect that the performance gains are higher for OpenCLIP because its attention maps naturally have more artifacts. Test-time registers also show slight boosts over the retrained DINOv2 with registers on mean IOU and mAP, suggesting that test-time registers lead to attention maps as clean as those from trained registers.
Mean IOU Pixel Accuracy mAP
DINOv2 38.3 87.6 80.6
DINOv2 + trained registers 33.9 91.4 79.0
DINOv2 + test-time registers 38.9 87.6 81.1
OpenCLIP 34.7 76.0 79.2
OpenCLIP + test-time registers 40.0 74.2 82.1
Table 4: Zero-shot segmentation results on ImageNet. We use the mean CLS attention maps from the last layer and find that test-time registers outperform the original models and DINOv2 models with trained registers on mean IOU and mAP with small drops in pixel accuracy.
@@ 5.3 Unsupervised object discovery
Test-time registers improve unsupervised object discovery. We report correct localization (corloc) scores in Table 5. We find that LOST performance improves significantly over the base model on features computed with our method for DINOv2 (increase of ${\sim}{21}$ corloc). Adding the test-time register closes the gap between the baseline model and a model with trained registers, reaching within ${\sim}{0{-}2}$ corloc. This suggests that the test-time register mimics the role of trained registers on this task. However, we note that test-time registers only marginally affect the results for OpenCLIP, a phenomenon also observed in Darcet et al. (2024) with trained registers. Further analysis can be found in Section A.9.
VOC 2007 VOC 2012 COCO 20k
DINOv2 32.2 36.8 25.4
DINOv2 + trained registers 56.2 60.2 42.3
DINOv2 + test-time registers 53.8 57.9 41.9
OpenCLIP 30.8 35.9 23.4
OpenCLIP + test-time registers 30.9 35.9 23.6
Table 5: Unsupervised Object Discovery with LOST (Siméoni et al., 2021). We report the best correct localization score (corloc) over the key, query, and value features from the last four layers. Adding test-time registers significantly boosts performance for DINOv2, effectively closing the gap with DINOv2 trained with registers.
@@ 5.4 Applying test-time registers to vision-language models
[…]
@@
Test-time registers maintain performance. We report our benchmark results in Table 6. Overall, we observe that adding a test-time register preserves performance for multi-modal processing. We do not pass the register to the LLM; we leave exploring its use as global memory for the LLM or leveraging multiple registers for adaptive computation to future work.
Test-time registers improve VLM interpretability. While adding a test-time register has only a minor impact on performance, it improves the interpretability of attention maps, as illustrated in Figure 7. We visualize the norms of the patch outputs from the vision encoder, highlighting the presence of outlier tokens. Next, we visualize the average attention – aggregated across all layers and heads of the language model – from the token responsible for answering the question to the visual tokens. Without registers, the outlier visual tokens create artifacts in the language model’s attention. However, adding a test-time register removes outliers and results in more interpretable cross-modal attention. This provides clearer insights into the model’s behavior (e.g., the attention map shows incorrect localization of the “man closest to us"). We provide more visualizations in Section A.10. Our results suggest that test-time registers can support diagnostics and the development of more transparent multimodal systems.
@@
5.5 Typographic attacks
We demonstrate here additional evidence of our ability to shift high-norm artifacts arising from register neurons. Azuma & Matsui (2023) showed that CLIP is vulnerable to typographic attacks, where images can be misclassified based on written text in the image rather than the depicted object (Figure 8). We strategically place high-norm artifacts within an image to mask out adversarial patches while preserving semantic content. This intervention is highly localized and relies on activating only a small fraction of neurons to produce targeted changes in the image representation.
Experimental setup. Following Azuma & Matsui (2023), we use OpenAI CLIP (Radford et al., 2021) and identify register neurons with Algorithm 1 on this model. Then, we localize the largest region of text in the image with OCR and, using the register neurons, move high-norm outliers onto the patches corresponding to the text location. We evaluate on the RTA-100 dataset (Azuma & Matsui, 2023), which contains approximately 1000 images with text written on top of an object from one of 100 classes. We calculate zero-shot accuracy by comparing CLIP’s image embedding to the “real” and “attack” labels’ text embedding.
Model Attack success %
CLIP 50.5
w/ pixel ablation 7.6
w/ neuron ablation 7.5
Results. Qualitative results in Figure 8 show that shifting the activation maps of register neurons early in the model’s computation causes outlier patches to later appear corresponding to the text area. Table 7 presents the attack success rate and find that it significantly drops after our intervention, indicating that our method masks part of the model’s internal representation without harming semantic content. Our results suggest that patch information at outliers is lost from the model computation rather than transferred elsewhere. We also present results of an alternative masking procedure in the input space, setting areas corresponding to text to the mean pixel value of the region. Our intervention matches this method’s performance using only a sparse modification – repurposing CLIP’s internal mechanisms by modifying roughly $0.02\%$ of all neurons, compared to masking ${\sim}10\%$ of the input. This highlights the influence of register neurons in guiding CLIP’s visual representation and output behavior.
@@ 6 Discussion, limitations, and future work
While our analysis shows that we can steer the location of high-norm tokens, we only addressed one component type that is responsible for their creation – neurons, while neglecting other possible elements that can contribute to their formation, such as attention layers or positional encodings (Yang et al., 2024). The edited tokens result in slight performance differences with their learned counterpart (Table 1), suggesting that the test-time registers are not fully equivalent to the high-norm tokens. Finally, similarly to Darcet et al. (2024), we mostly focus on individual models (CLIP-ViT-B/16 and DINOv2-ViT-L/14). We do not present results on other, less commonly used, pretrained ViTs, and leave it for future work.
The mechanism that we found points to an intriguing property about model neurons – not all neurons have a feature-related role. Register neurons, for example, are responsible for igniting high-norm tokens – an image-independent role, and can not be discovered by correlating their activations to the image features. Uncovering other similar image-independent roles of neurons can shed additional light on the computational process of deep neural networks. We plan to develop a methodology for such automatic discovery in future work.
@@ 7 Acknowledgments
We thank Katie Luo, Lisa Dunlap, Yaniv Nikankin, and Arjun Patrawala for their feedback on our paper. AD is supported by the US Department of Energy Computational Science Graduate Fellowship. YG is supported by the Google Fellowship. Additional support came from the ONR MURI grant.

### v5

@@ 5 Experiments
We evaluate how adding test-time registers affects the downstream performance of models trained without registers. We begin by detailing the evaluated models, then compare performance across classification, dense prediction, unsupervised object discovery, and zero-shot segmentation tasks, finding that test-time registers perform comparably to their retrained counterparts. Finally, we apply test-time registers to an off-the-shelf vision-language model to improve its interpretability. We present additional experiments on preventing typographic attacks in Section˜A.8.
Models. We evaluate using DINOv2 (Oquab et al., 2024) and OpenCLIP (Cherti et al., 2023). For DINOv2, we use the publicly released ViT-L/14 checkpoints trained on LVD-142M, including both the standard model and a variant trained with four registers. These two models serve as our baselines, while our approach applies edits to the standard model. For OpenCLIP, we evaluate the ViT-L/14 and ViT-B/16 models trained on LAION-2B (Schuhmann et al., 2022). As no checkpoints with trained registers are available for OpenCLIP, we only compare our approach to the standard models. We present results on larger models in Section˜A.7.
5.1 Classification and dense prediction
The results in Table˜3 show that models edited with test-time registers maintain their performance on ImageNet classification with slight performance gains over the base model on segmentation and depth estimation tasks. These improvements are consistent with the performance gains observed in DINOv2 explicitly trained with registers. Thus, we demonstrate that intervening on a model with test-time registers does not degrade the model’s representations and, in some cases, even enhances performance on prediction tasks. While we observe a small improvement in classification performance with the DINOv2 model trained with registers, we note that this was an independently trained model. Hence, the difference could be due to variations in initialization and training dynamics, rather than the effect of trained registers fixing artifacts. We report $95\%$ confidence intervals for these results in Section˜A.13 and find that performance differences remain consistent.
@@ 5.1 Classification and dense prediction
DINOv2 ViT-L/14 86.4 48.3 0.388
w/ trained registers 86.7 49.1 0.382
w/ test-time register 86.4 49.1 0.378
OpenCLIP ViT-B/16 77.4 40.1 0.603
w/ test-time register 77.5 40.3 0.596
Table 2: Linear probing results. The performance of models with test-time registers maintains or improves performance over the original models, and largely matches models with trained registers.
@@
w/ test-time register 76.4
@@
w/ test-time register 71.3
@@
Zero-shot classification. We evaluate zero-shot ImageNet classification with OpenCLIP to assess whether test-time registers preserve the semantic structure of the original representation space. Unlike linear probing, where a trained linear head can compensate for small shifts in representation, zero-shot classification is more sensitive to such changes. We compare zero-shot performance before and after applying test-time registers across both ViT-L/14 and ViT-B/16 in Table˜3, observing that the intervention does not sacrifice performance.
@@ 5.2 Zero-shot segmentation
To validate that test-time registers result in more interpretable attention maps, we follow a standard protocol for evaluating heatmap-based explainability methods (Hooker et al., 2019) – binarizing the heatmap into a foreground/background segmentation map, and evaluating its segmentation quality. We compute the mean attention map for the [CLS] token in the last layer (Chefer et al., 2021) for the original model without registers, the model with test-time registers, and a model trained with registers if available. We evaluate zero-shot segmentation performance on ImageNet-segmentation (Guillaumin et al., 2014), which contains 4,276 images from the ImageNet validation set with annotated segmentations.
Results. We present our zero-shot segmentation scores in Table˜5. We also qualitatively compare the attention maps for DINOv2 with test-time and trained registers in Figure˜5, demonstrating similarly high-quality attention maps. Using test-time registers on both DINOv2 and OpenCLIP outperforms the original model on mean IOU and mAP with minimal drops in pixel accuracy. Test-time registers also show slight boosts over the retrained DINOv2 with registers on mean IOU and mAP, suggesting that test-time registers lead to attention maps as clean as those from trained registers.
@@ 5.3 Unsupervised object discovery
Test-time registers improve unsupervised object discovery. We report correct localization (corloc) scores in Table˜5, using the best result across key, query, and value features from the final four layers. We find that LOST performance improves significantly over the base model on features computed with our method for DINOv2 (increase of ${\sim}{21}$ corloc). Adding the test-time register closes the gap between the baseline model and a model with trained registers, reaching within ${\sim}{0{-}2}$ corloc. This suggests that the test-time register mimics the role of trained registers on this task. However, we note that test-time registers only marginally affect the results for OpenCLIP, a phenomenon also observed in Darcet et al. (2024) with trained registers. Further analysis can be found in Section˜A.14.
mIoU Pix. Acc. mAP
DINOv2 ViT-L/14 $38.3$ $87.6$ $80.6$
w/ trained registers $33.9$ $91.4$ $79.0$
w/ test-time register $38.9$ $87.6$ $81.1$
OpenCLIP ViT-B/14 $34.7$ $76.0$ $79.2$
w/ test-time register $40.0$ $74.2$ $82.1$
Table 4: Zero-shot segmentation results on ImageNet. We use the last layer’s mean [CLS] attention maps and find that test-time registers outperform the original models and DINOv2 models with trained registers on mean IOU and mAP.
VOC07 VOC12 COCO
DINOv2 ViT-L/14 $32.2$ $36.8$ $25.4$
w/ trained registers $56.2$ $60.2$ $42.3$
w/ test-time register $53.8$ $57.9$ $41.9$
OpenCLIP ViT-B/16 $30.8$ $35.9$ $23.4$
w/ test-time register $30.9$ $35.9$ $23.6$
Table 5: Unsupervised Object Discovery with LOST (Siméoni et al., 2021). Adding test-time registers significantly boosts performance for DINOv2, effectively closing the gap with DINOv2 trained with registers.
@@ 5.4 Applying test-time registers to vision-language models
Evaluation setting. Using Algorithm˜1, we find a set of 100 register neurons (out of $\sim$}100K neurons) from CLIP ViT-L/14 vision encoder of LLaVA-Llama-3-8B. [^3] https://huggingface.co/xtuner/llava-llama-3-8b. We then create a test-time register to collect their activations and evaluate the model on the eight main benchmarks from the VLMEvalKit toolkit (Duan et al., 2024). The benchmarks span across OCR, chart interpretation, visual Q/A, etc.
@@
Test-time registers maintain performance. We report our benchmark results in Table˜6. Overall, we observe that adding a test-time register preserves performance for multi-modal processing. We do not pass the register to the LLM; we leave exploring its use as global memory for the LLM or leveraging multiple registers for adaptive computation to future work.
Test-time registers improve VLM interpretability. While adding a test-time register has only a minor impact on performance, it improves the interpretability of cross-modal attention maps, as illustrated in Figure˜6. We visualize the norms of the patch outputs from the vision encoder, highlighting the presence of outlier tokens. Next, we visualize the average attention – aggregated across all layers and heads of the language model – from the token responsible for answering the question to the visual tokens. Without registers, the outlier visual tokens create artifacts in the language model’s attention. However, adding a test-time register removes outliers and results in more interpretable text-to-vision attribution. This provides clearer insights into the model’s behavior (e.g., the attention map shows incorrect localization of the “man closest to us"). We provide more visualizations in Section˜A.15.
@@
6 A mathematical model for register neurons
In this section, we present a simple mathematical model describing how register neurons can induce high-norm tokens that attract excess attention. We then relate this behavior to our empirical findings. Prior work has examined the emergence of outlier tokens in large language models and proposed several explanations, most notably the no-op (no-operation) hypothesis (Clark et al., 2019; Kobayashi et al., 2020; Bondarenko et al., 2023), which attributes these effects to the softmax constraint enforcing that attention weights sum to one. As a result, the attention operation forces aggregating some information from other tokens, even when current embeddings already contain sufficient information for the task. This constraint can direct excess attention towards low-information tokens, creating attention sinks (Xiao et al., 2024). We show that, under certain conditions, register neurons can produce high-norm tokens that give rise to such attention sinks within the no-op framework.
6.1 Analytical model
Set-Up. We consider the MLP layer from the penultimate Transformer block followed by the full final block (attention + MLP) of a Vision Transformer without normalization layers. Let the input sequence to the penultimate MLP be a task-specific [CLS] token and $n$ patch tokens: $x^{(0)}_{cls},x^{(0)}_{1},...x^{(0)}_{n}\in\mathbb{R}^{d}$ The MLPs will consist of two linear layers (no bias) and an activation function $\phi$ (e.g., ReLU or GELU). We denote the weight matrices of the penultimate block’s MLP as $W_{\text{in}}^{(1)}\in\mathbb{R}^{d\times d_{\text{mlp}}},W_{\text{out}}^{(1)}\in\mathbb{R}^{d_{\text{mlp}}\times d}$, and those of the final block’s MLP as $W_{\text{in}}^{(2)}\in\mathbb{R}^{d\times d_{\text{mlp}}},W_{\text{out}}^{(2)}\in\mathbb{R}^{d_{\text{mlp}}\times d}$. There will be one attention head with identity projection matrices: $W_{Q}=W_{K}=W_{V}=W_{O}=I_{d}\in\mathbb{R}^{d\times d}$. Finally, the output task-specific head is a linear projection that uses the [CLS] token’s embedding for prediction: $W_{\text{head}}\in\mathbb{R}^{d\times c}$, where $c$ is the number of output classes (or regression targets). Our analysis is carried out in the regime where both $W_{\text{in}}^{(2)}$, $W_{\text{head}}$ are low-rank, i.e., $\mathrm{rank}(W_{\text{in}}^{(2)})<\mathrm{min}(d,d_{\text{mlp}})$ and $\mathrm{rank}(W_{\text{head}})<\mathrm{rank}(d,c)$. Hence, both matrices possess non-trivial kernels.
Proposition 1 (Register neuron induces attention sink and no-op attention)
Let $u_{1}=(W_{\text{in}}^{(1)})_{:,1}$ be a register neuron and $u_{2}=(W_{\text{out}}^{(1)})_{1,:}$ be the corresponding row in the MLP’s second weight matrix, with $\|u_{2}\|\gg\|(W_{\text{out}}^{(1)})_{j,:}\|$ for $j\neq 1$. If both $u_{1},u_{2}\in\ker(W_{\text{in}}^{(2)\top})\cap\ker(W^{\top}_{\text{head}})$ and a token $x_{s}$ is aligned with $u_{1}$, i.e., $x^{(0)}_{s}=\beta u_{1}$ for some $\beta>0$, then $x_{s}$ becomes an attention sink and the attention layer has no contribution to the model’s output.
Proof. See Section˜A.1.
Corollary 1 (Register neuron induces implicit attention bias)
The contribution of register neurons is equivalent to the attention mechanism with explicit bias terms ($\mathbf{k}^{\prime},\mathbf{v}^{\prime}$) as in Sun et al. (2024):
$\text{Attention}(Q,K,V;\mathbf{k}^{\prime},\mathbf{v}^{\prime})=\text{softmax}\left(Q\begin{bmatrix}K^{T}&\mathbf{k}^{\prime}\end{bmatrix}\right)\begin{bmatrix}V\\ \mathbf{v}^{\prime T}\end{bmatrix}$
Proof. See Section˜A.1.
6.2 Empirical validation on OpenCLIP
This mathematical model captures three distinct observed behaviors of register neurons and high norm tokens. We validate this correspondence qualitatively and quantitatively. Notably, our framework admits an equivalent attention formulation augmented with explicit bias terms – values that can be derived directly from register neurons without any additional training.
(1) Task-Irrelevant Attention Sinks: The token that achieves the attention sink behavior in our mathematical model lives in the kernel, or null space, of the task specific head, indicating it is not informative of the task. This aligns with the observation that register neurons fire on seemingly random, uninformative regions such as background or uniform texture (e.g., see Figure˜3). In the case of language models, attention sinks have been observed to appear in semantically low-information tokens such as <BOS> or delimiters (Xiao et al., 2024).
(2) Lack of Sink Ruins Performance: Our framework suggests that zeroing out the activation of a register neuron would disrupt the “no-op” attention behavior, thus unnecessarily altering the [CLS] token from its intended representation and substantially degrading model performance. To test this, we zero-out the 10 register neurons in OpenCLIP ViT-B/16 and observe a drop in IN1k zero-shot performance from 70.4% to 55.6%. As a baseline, zeroing out a random set of 10 neurons over three trials results in minimal change: $69.3\%\pm 1.1$.
(3) Register Neurons Induce an Implicit Attention Bias:  Corollary˜1 aligns with prior observations (Sun et al., 2024; Gu et al., 2024) that high-norm tokens induce an implicit bias term in the attention mechanism. Sun et al. (2024) proposed learning biases in the attention layers of large language models to remove outliers. Specifically, given the query, key, and value features from $T$ tokens $Q,K,V\in\mathbb{R}^{T\times d}$ for each attention head, they train additional parameters $\mathbf{k}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{d}$ in:
$\text{Attention}(Q,K,V;\mathbf{k}^{\prime},\mathbf{v}^{\prime})=\text{softmax}\left(\frac{Q\begin{bmatrix}K^{T}&\mathbf{k}^{\prime}\end{bmatrix}}{\sqrt{d}}\right)\begin{bmatrix}V\\ \mathbf{v}^{\prime T}\end{bmatrix}$ (1)
To test that register neurons implicitly induce attention biases as predicted by Corollary˜1, we use OpenCLIP ViT-B/16. For each attention head, $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ are set to the mean key and value vectors of the test-time register token, which captures the aggregate contribution of the register neurons, averaged across 1000 images. Instead of a test-time register, we zero out the register neurons and inject the constant vectors $\mathbf{k}^{\prime}$ and $\mathbf{v}^{\prime}$ directly into the attention computation as above at interference. This maintains the $71.3\%$ IN1k zero-shot performance while suppressing artifacts in attention maps (Figure˜7) and mitigating outliers completely from the residual stream (details in Section˜A.2).
Figure 7: Register-based attention bias mitigates artifacts. Mean [CLS] attention maps from the last layer of OpenCLIP ViT-B/16 show that attention bias terms derived from test-time registers (Equation˜1) suppress artifacts, matching the effect of explicitly using test-time registers.
@@ 7 Discussion, limitations, and future work
While our analysis shows that we can steer the location of high-norm tokens, we only addressed one component type that is responsible for their creation – neurons, while neglecting other possible elements that can contribute to their formation, such as attention layers or positional encodings (Yang et al., 2024). The edited tokens result in slight performance differences with their learned counterpart (Table˜1), suggesting that the test-time registers are not fully equivalent to the high-norm tokens. Finally, similarly to Darcet et al. (2024), we mostly focus on individual models (CLIP, DINOv2). We do not present results on other, less commonly used, pretrained ViTs, and leave it for future work.
The mechanism that we found points to an intriguing property about model neurons – not all neurons have a feature-related role. Register neurons, for example, are responsible for igniting high-norm tokens – an image-independent role – and cannot be discovered by correlating their activations to the image features. Similar high-norm tokens have also been observed in language models (Xiao et al., 2024; Dettmers et al., 2022), suggesting that large language models may rely on a related mechanism. Uncovering other similar input-independent roles of neurons can shed additional light on the computational process of deep neural networks. We plan to develop a methodology for such automatic discovery in future work.
@@ 8 Acknowledgments
We thank Katie Luo, Lisa Dunlap, Yaniv Nikankin, and Arjun Patrawala for their feedback on our paper. AD is supported by the US Department of Energy Computational Science Graduate Fellowship. YG is supported by the Google Fellowship. Additional support came from the ONR MURI grant and NSF IIS-2403305.